OpenStudy (anonymous):

Hello, Im trying to find Pr(|X|>15) where X is approximately a normal distribution. But my answer doesnt match the correct answer. I have no idea whats wrong.

OpenStudy (anonymous):

ok.. how did you solved it? once check 1-(your answer) matches with correct answer

OpenStudy (anonymous):

I rearranged it as 2-2(Pr(|X|>15)) this may be a mistake but I dont see how this rearrangement is wrong. Afterwards I just found the z-score and looked it up on a table.

OpenStudy (anonymous):

sorry I mean 2-2Pr(|X|<=15)

OpenStudy (anonymous):

I think just 2P(X<15) should do it.

OpenStudy (anonymous):

Do you have the mean and standard deviation of X?

OpenStudy (anonymous):

@CliffSedge not x<15, it should be x>15

OpenStudy (anonymous):

Ah, yes, I see the error in considering the ||

OpenStudy (anonymous):

Though, I think because of symmetry, it doesn't matter.

OpenStudy (anonymous):

@witcon .. see pr(|x|>15)= pr(x<-15) +pr(x>15) or 1- pr(-15<x<15) since pr(x<-15) = pr(x>15) 2*pr(x>15) will do, @CliffSedge since it is normal dist. mean = 0, SD =1

OpenStudy (anonymous):

It's only N(0,1) if standard normal.

OpenStudy (anonymous):

Its a normal dist. not normal standard, I have to convert using z-score

OpenStudy (anonymous):

u=20 var=25

OpenStudy (anonymous):

Also theloser's derivation gives the wrong answer

OpenStudy (anonymous):

calculate this 2*pr(x<-15)

OpenStudy (anonymous):

0.3174, the correct answer is 0.8413

OpenStudy (anonymous):

Is there something else |X| could mean? Otherwise I dont see where Im going wrong.

OpenStudy (phi):

mu= 20 and sigma = 5 Pr(x>15) is 0.8413 (see http://davidmlane.com/hyperstat/z_table.html) note: you might have to refresh the page before it takes input Pr(x<-15) is way off in the left tail, essentially 0

OpenStudy (anonymous):

yes youre right Pr(X>15) is 0.8413 but thats not the same thing as Pr(|X|>15) is it?

OpenStudy (phi):

you want the sum of Pr(x>15) + Pr(x<-15) but -15 is 7 std dev below the mean at 20

OpenStudy (phi):

you change |x|>15 to 2 relations +x > 15 and -x > 15 --> x< -15

OpenStudy (anonymous):

wait sorry Pr(X>15) is 1- Pr(X< 15) which is 1- 0.8413 or 0.1587

OpenStudy (phi):

? 20 is the mean. the probability of being 20 or higher is 0.5 (half the samples are 20 or greater) the prob of 15 or greater is (obviously?) greater than 0.5

OpenStudy (anonymous):

am I doing the z-score wrong, (15-20)/5 should be -1, meaning Z < -1 or Z > 1 or 1-(Z<1) which is 0.1587

OpenStudy (anonymous):

Im making a mistake somewhere there

OpenStudy (phi):